Two oscillations along the same line are
described by the equations
x1= 0.03 cos (10π t)
and x2= 0.03 cos (12π t)
Obtain the equation describing the resultant
motion due to superposition of these two
oscillations.
1
Expert's answer
2019-05-23T09:44:08-0400

The equation describing the resultant motion is given by formula


"x=0.03 (\cos 10π t + \cos 12π t) (1)"

Using (1) we got:


"x=0.06 (\cos 11π t \cos -π t)= 0.06 (\cos 11π t \cos π t)"


Answer:




"0.06 (\cos 11π t \cos π t)"